7c2-3c-4=0 by completing the squares
Answer by lwsshak3(11628)

7c2-3c-4=0 by completing the squares
7c^2-3c-4=0
complete the square:
7(c^2-(3/7)c+(9/196)-9/28-4=0
7(c-(3/14)^2-121/28=0
divide by 7
(c-(3/14)^2=121/196
take sqrt of both sides
c-3/14=±11/14)
c=3/14±11/14
c=(3/14+11/14=14/14=1
or
c=(3/14-11/14=-8/14≈-0.571